# FPGA-tailored algorithms for real-time decoding of quantum LDPC codes

### v1

@@ 1 Introduction and summary of main results
Our first technical contribution is an FPGA-tailored variant, filtered-OSD, which avoids explicit matrix inversion by solving $H_{S^{\prime}}x=\sigma$ directly on a smaller, potentially rank-deficient submatrix $H_{S^{\prime}}$ [panteleev2021degenerate]. We obtain $S^{\prime}$ from a pre-decoder by discarding low-probability columns. Because $|S^{\prime}|\ll\operatorname{rank}(H)$, column sorting and linear solving are both substantially accelerated. Sparsity further reduces cost because many rows of $H_{S^{\prime}}$ are zero and can be skipped by the solver. This approach requires an FPGA-tailored solver for rank-deficient binary systems, described next.
@@
We provide a new linear-time systolic-array algorithm that solves $Ax=b$ for any matrix $A\in\mathbb{F}_{2}^{m\times n}$ and vector $b\in\mathbb{F}_{2}^{m}$. Systolic arrays are a class of highly parallel algorithms that implement computations on grids of processing elements and pass data locally each cycle, making them particularly well suited to FPGAs. Existing FPGA solvers typically assume that $A$ has full rank [gentleman1982matrix, Hochet1989, cosnard1986matching, robert1985resolution, wang1993systolic, parkinson1984compact, Rupp2006, rupp2011hardware, jasinski2010gf2, scholl2013hardware, scholl2014hardware, Shoufan2010, wang2017fpga, bascones2025exploring], while the only prior solver for arbitrary $A$ [hu2024universal] employs relatively complex processing elements. Our solver is simpler and may be of independent interest; it forms the computational core of our filtered-OSD and cluster decoder.
@@
This overall conclusion appears unlikely to change without substantial new algorithmic advances in OSD or cluster decoding algorithms. Although our cycle counts and resource estimates are obtained at the architectural level rather than from full Verilog implementations (see Section LABEL:sec:results), the observed performance gaps are simply too large to be bridged by modest clock-frequency variations. Here is a clearer and slightly tighter version with no added length: OSD could improve logical error rates by adding a combination-sweep (CS) stage [roffe2020], but not enough to rival Relay [muller2025improved]. Moreover, although CS adds negligible overhead on CPUs (where OSD’s elimination cost dominates over the matrix–vector multiplications in CS) it would cause a substantial slowdown on FPGAs, where elimination and matrix–vector multiplication have comparable linear parallel-time cost. For similar reasons, although clustering could reuse partial elimination results [hillmann2024localized] for a speedup on CPUs, we expect little advantage on FPGAs where elimination is less of a bottleneck. Cluster decoders might also benefit from non-uniform, pre-decoder-biased growth [wolanski2024ambiguity], but such extensions would introduce conditional logic and data-management overheads that reduce throughput.

### v2

@@ 1 Introduction and summary of main results
Our first technical contribution is an FPGA-tailored variant, filtered-OSD, which avoids explicit matrix inversion by solving $H_{S^{\prime}}x=\sigma$ directly on a smaller, potentially rank-deficient submatrix $H_{S^{\prime}}$ (similar to Alg.2 in Ref [panteleev2021degenerate]). We obtain $S^{\prime}$ from a pre-decoder by discarding low-probability columns. Because $|S^{\prime}|\ll\operatorname{rank}(H)$, column sorting and linear solving are both substantially accelerated. Sparsity further reduces cost because many rows of $H_{S^{\prime}}$ are zero and can be skipped by the solver. This approach requires an FPGA-tailored solver for rank-deficient binary systems, described next.
@@
We provide a new linear-time systolic-array algorithm that solves $Ax=b$ for any matrix $A\in\mathbb{F}_{2}^{m\times n}$ and vector $b\in\mathbb{F}_{2}^{m}$. Systolic arrays are a class of highly parallel algorithms that implement computations on grids of processing elements and pass data locally each cycle, making them particularly well suited to FPGAs. Existing FPGA solvers typically require $A$ has either full column or row rank (or both) [gentleman1982matrix, Hochet1989, cosnard1986matching, robert1985resolution, wang1993systolic, parkinson1984compact, Rupp2006, rupp2011hardware, jasinski2010gf2, scholl2013hardware, scholl2014hardware, Shoufan2010, wang2017fpga, bascones2025exploring], while the only prior solver for arbitrary $A$ [hu2024universal] employs relatively complex processing elements. Our solver is simpler and may be of independent interest; it forms the computational core of our filtered-OSD and cluster decoder.
@@
This overall conclusion appears unlikely to change without substantial new algorithmic advances in OSD or cluster decoding algorithms. Although our cycle counts and resource estimates are obtained at the architectural level rather than from full Verilog implementations (see Section LABEL:sec:results), the observed performance gaps are simply too large to be bridged by modest clock-frequency variations. OSD could improve logical error rates by adding a combination-sweep (CS) stage [roffe2020], but not enough to rival Relay [muller2025improved]. Moreover, although CS adds negligible overhead on CPUs (where OSD’s elimination cost dominates over the matrix–vector multiplications in CS) it would cause a substantial slowdown on FPGAs, where elimination and matrix–vector multiplication have comparable linear parallel-time cost. For similar reasons, although clustering could reuse partial elimination results [hillmann2024localized] for a speedup on CPUs, we expect little advantage on FPGAs where elimination is less of a bottleneck. Cluster decoders might also benefit from non-uniform, pre-decoder-biased growth [wolanski2024ambiguity], but such extensions would introduce conditional logic and data-management overheads that reduce throughput.